OpenStudy (anonymous):

Please explain to me the steps in solving 2.14 x - 12.18 = -5.76

OpenStudy (anonymous):

You need to isolate the x

OpenStudy (anonymous):

you must find the x

OpenStudy (anonymous):

First move - 12.8 to the other side

OpenStudy (anonymous):

So add 12.8 to both sides of equation

OpenStudy (anonymous):

Then divide by 2.14

OpenStudy (anonymous):

Divide by 2.14 on both sides.

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

I'm sorry, I don't get it, I'm really bad at math :/ I never know where to put numbers and whatnot... would you mind explaining it to me again?

OpenStudy (anonymous):

Yes, no problem. You have to isolate X from all the other numbers, and put them on the other side of the equation.

OpenStudy (anonymous):

So since - 12.8 is on the X side, you have to add 12.8 to both sides of the equation so that it cancels out on one side and it's added on the other side. Remember whatever you do to one side of the equation, you also have to do it to the other side.

OpenStudy (anonymous):

Then since 2.14 is being multiplied to X just divide both sides by 2.14.

OpenStudy (anonymous):

Okay, and what do you mean by "both sides of the equation?" Is it 12.18 + 2.14, or 12.18 + -5.76?

OpenStudy (anonymous):

like this: \[2.4x-12.8+12.8=-5.76+12.8\]

OpenStudy (anonymous):

then divide by 2.4 on both sides so that X can be isolated.

OpenStudy (anonymous):

x=-2.933333333

OpenStudy (anonymous):

sorry I messed up, it's 2.14, not 2.4!

OpenStudy (anonymous):

and 7.04 is a positive number too! I was on my phone trying to answer!

OpenStudy (anonymous):

\[\frac{ 2.14x }{ 2.14 }=\frac{ 7.04 }{ 2.14 } \]

OpenStudy (anonymous):

x=3.289719626

OpenStudy (anonymous):

to check your answer, just substitute the value of X into the original equation:

OpenStudy (anonymous):

2(3.289719626)-12.8=-5.76

OpenStudy (anonymous):

Ohhh, okay, so the answer is 3, correct?

OpenStudy (anonymous):

3.289719626

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

You're welcome!